# DC-VSR: Spatially and Temporally Consistent Video Super-Resolution with Video Diffusion Prior

### v1

@@ Abstract.
Video super-resolution (VSR) aims to reconstruct a high-resolution (HR) video from a low-resolution (LR) counterpart. Achieving successful VSR requires producing realistic HR details and ensuring both spatial and temporal consistency. To restore realistic details, diffusion-based VSR approaches have recently been proposed. However, the inherent randomness of diffusion, combined with their tile-based approach, often leads to spatio-temporal inconsistencies. In this paper, we propose DC-VSR, a novel VSR approach to produce spatially and temporally consistent VSR results with realistic textures. To achieve spatial and temporal consistency, DC-VSR adopts a novel Spatial Attention Propagation (SAP) scheme and a Temporal Attention Propagation (TAP) scheme that propagate information across spatio-temporal tiles based on the self-attention mechanism. To enhance high-frequency details, we also introduce Detail-Suppression Self-Attention Guidance (DSSAG), a novel diffusion guidance scheme. Comprehensive experiments demonstrate that DC-VSR achieves spatially and temporally consistent, high-quality VSR results, outperforming previous approaches.
Figure 1. Real-world video super-resolution results of our proposed methods and state-of-the-art models: RealViformer (Zhang and Yao, 2024), Upscale-A-Video (Zhou et al., 2024), MGLD (Yang et al., 2024a). DC-VSR shows outstanding super-resolution performance, in respect of not only quality but also spatial and temporal consistency.
@@ 1. Introduction
To achieve successful VSR, it is essential to generate realistic HR details while maintaining both spatial and temporal consistency. To this end, over the past few decades, numerous approaches have been proposed, ranging from traditional methods such as interpolation techniques and model-based optimization approaches to recent neural network-based ones (Zhang and Yao, 2024). Nevertheless, most of these approaches struggle to produce realistic HR details due to a lack of effective priors that model high-frequency details as shown in Fig. 1(b).
Recently, diffusion models that provide powerful generative priors for natural images, have been exploited for VSR to achieve high-quality results with realistic textures. Yang et al. (2024a) and Zhou et al. (2024) utilize image diffusion models as their generative priors and successfully restore detail-rich HR videos from real-world LR videos. However, adopting image diffusion models poses challenges in maintaining spatial and temporal consistency due to the design of image diffusion models, which typically target single images with limited spatial sizes due to significant memory usage. To achieve temporally-consistent VSR using image diffusion models, these approaches adopt additional temporal layers, frame information propagation with motion compensation, and overlapping of temporal windows. Furthermore, to handle large video frames, frames are split into overlapping tiles, processed individually, and then merged. Despite these efforts, the resulting videos often suffer from spatially and temporally inconsistent results due to the inherent randomness of image diffusion models and their tile-based approach (Fig. 1(c) and (d)).
In this paper, we propose DC-VSR (Diffusion-based Consistent VSR), a novel VSR approach that, for the first time, leverages a video diffusion prior. Our approach produces spatially and temporally consistent results with realistic textures, given an LR video of arbitrary length and spatial size (Fig. 1(e)). We leverage Stable Video Diffusion (SVD) (StabilityAI, 2023) as a video diffusion prior, which provides powerful generative capabilities for restoring high-quality and temporally-consistent details. However, exploiting an existing video diffusion model for consistent VSR over long video clips with large frames is not straightforward, since existing models are designed to synthesize a limited number of small frames, similar to image diffusion models.
@@
  • We propose a Spatial Attention Propagation (SAP) scheme that injects subsampled features representing the entire area of a video frame to different tiles, ensuring spatial consistency.
  • We propose a Temporal Attention Propagation (TAP) scheme for sharing information across temporally distant frames, achieving temporal consistency.
  • We introduce Detail-Suppression Self-Attention Guidance (DSSAG), which enhances the quality of synthesized video frames without any additional computational overhead.
@@ 2. Related Work
VSR methods can broadly be classified into two main categories based on the use of generative prior. VSR models without generative prior (Youk et al., 2024; Chan et al., 2021, 2022a, 2022b; Xie et al., 2023; Zhang and Yao, 2024) are typically trained using reconstruction losses, such as the mean squared error (MSE), augmented with additional techniques designed to enhance temporal consistency. Although they restore HR videos closely matching the input LR videos, they fail to generate complex details, producing blurry results due to the significant ill-posed nature of the problem. To achieve VSR with rich details, generative prior-based VSR methods (Rota et al., 2024; Xu et al., 2024; Chen et al., 2024; Zhou et al., 2024; Yang et al., 2024a) have recently emerged. These methods fine-tune GANs (Goodfellow et al., 2014) or diffusion-based image generation models by modifying them to take LR inputs as conditions and incorporating additional modules to consider temporal consistency. However, maintaining consistency across frames with detailed textures is still challenging (Blattmann et al., 2023a, b), leading to temporal flickering artifacts. To address this challenge, we propose to leverage video generative priors instead of image generative priors for the first time. By fine-tuning a high-quality video diffusion model (StabilityAI, 2023), we successfully achieve not only rich details but also improved temporal consistency in a simple yet effective manner.
@@
As deep learning techniques have advanced, various image super-resolution (ISR) methods have been proposed. SRCNN (Dong et al., 2015) was the first to introduce a CNN-based SR method. Subsequent studies have applied various network architectures, such as residual networks (Kim et al., 2016a; Ledig et al., 2017), dense networks (Zhang et al., 2018b; Wang et al., 2018b, 2021), Laplacian pyramid networks (Lai et al., 2017), back-projection networks (Haris et al., 2018), recursive structures (Kim et al., 2016b; Tai et al., 2017), and channel attention (Zhang et al., 2018a). While these methods have demonstrated promising SR results, they often struggle to recover high-frequency details. To address this, approaches leveraging the generative prior of pretrained generative models have been introduced exploiting GANs (Chan et al., 2020; Menon et al., 2020) or Diffusion models (Wang et al., 2024; Yu et al., 2024; Lin et al., 2024; Yang et al., 2024b; Wu et al., 2024b; Sun et al., 2024; Wu et al., 2024a). Since the application of these methods to videos leads to severe flickering artifacts due to the absence of temporal consistency, VSR research has been active in addressing this issue.
@@
Diffusion models face two main challenges in generating higher spatial or temporal resolutions: quality degradation from discrepancies between training and inference conditions, and substantial computational costs. To address these challenges, tile-based generation approaches have been proposed (Bar-Tal et al., 2023; Lee et al., 2023; Du et al., 2024; Zhou et al., 2024; Yang et al., 2024a). These methods divide the input into overlapping tiles, perform diffusion sampling on each tile individually, and subsequently merge the tiles. However, their approaches do not consider inter-tile consistency, often generating inconsistent details for the same content across split tiles in an image or distant frames (Zhou et al., 2024; Yang et al., 2024a). Our proposed methods, SAP and TAP, propagate information across spatial and temporal tiles, producing spatially and temporally consistent VSR results.
@@ 3. DC-VSR
At each diffusion sampling step $t$, $x_{t}$ and $l$ are first concatenated in an interleaved manner, i.e., $[x_{t,1},l_{1},\cdots,x_{t,N},l_{N}]$, where $x_{t,i}$ is the noisy latent of the $i$-th HR video frame. The concatenated latents are then split into spatio-temporal tiles. We refer to each tile as $x^{\prime}_{t,m,n}$ where $m$ and $n$ are spatial and temporal indices. Each tile is processed through a denoising U-Net, and the processed tiles are merged to obtain the latent representation $x_{t-1}$ at the next sampling step $t-1$. We utilize spatio-temporal tiles of size $64\times 64\times 14$ in the latent space, corresponding to $512\times 512\times 14$ in the image space with a scaling factor of 8. In line with previous approaches (Zhou et al., 2024; Yang et al., 2024a), spatially and temporally neighboring tiles overlap by 50%. Overlapped tiles are blended in the tile merging step in our pipeline using gaussian blending. To achieve spatial and temporal consistency, DC-VSR alternatingly applies either SAP or TAP at each sampling step. This process is repeated until $t$ reaches 00. Finally, $x_{0}$ is fed to a decoder, producing the HR video $I^{\textrm{HR}}$.
DC-VSR employs a tile-based approach to handle lengthy videos with large frames with a video diffusion prior. However, naïvely splitting a video into tiles may introduce spatial and temporal inconsistencies. In image diffusion models like latent diffusion model (LDM) (Rombach et al., 2022), self-attention layers of the denoising U-Nets play a crucial role in ensuring spatial consistency within an image. Likewise, video diffusion models such as SVD (StabilityAI, 2023) leverage self-attention to achieve spatially and temporally coherent results. Specifically, the self-attention operation is defined as:
@@ 3.3. Detail-Suppression Self-Attention Guidance
SAG and PAG noticeably improve image synthesis quality, especially when combined with CFG. However, integrating them with CFG incurs substantial computational costs. Using $\epsilon_{\theta}(x_{t},c)$ instead of $\epsilon_{\theta}(x_{t})$ in Eqs. 5 and 6 allows this combination, but it necessitates running the denoising U-Net three times for SAG due to its blurring function. With PAG, the fixed level of perturbation complicates balancing the effects of CFG and PAG when combined. Consequently, PAG and CFG are typically applied separately, also leading to denoising U-Net three times.
Table 1. Quantitative comparisons of different VSR (x4) models on synthetic datasets (REDS4 and UDM10) and real datasets (VideoLQ). ISR, VSR and GP are image super-resolution tasks, video super-resolution tasks and generative prior. The best and second-best scores are written as bold and underline. $\text{DOVER}^{*}$ and $\text{tLP}^{*}$ denote DOVER ($\times$100) and tLP ($\times$100) respectively.
  Datasets Metrics ISR ISR with GP VSR VSR with GP
@@
Figure 3. Denoised results of unconditional term at the intermediate timestep (t=16 out of 25) (a) without and (b) with DSSAG.
@@ 4. Experiments
Implementation Details
To build DC-VSR, we fine-tune Image-to-Video Stable Video Diffusion (I2V-SVD) (StabilityAI, 2023), which adopts the LDM framework (Rombach et al., 2022) with the EDM (Karras et al., 2022) diffusion mechanism. We use the REDS dataset (Nah et al., 2019) to train our model. Following previous work (Chan et al., 2022b; Yang et al., 2024a), we merge 240 training videos and 30 test videos, reorganizing them into 266 training videos and 4 test videos, and refer to the latter as REDS4. We refer the reader to the supplementary material for more implementation details.
Evaluation Datasets
We use the REDS4 and UDM10 (Yi et al., 2019) datasets for VSR evaluation. We construct HR-LR video pairs from the datasets using the real-world degradation pipeline of Chan et al. (2022b), which applies random blur, resizing, noise, JPEG compression, and video compression. Additionally, we use the VideoLQ dataset (Chan et al., 2022b) as a real-world LR dataset.
@@ 4.1. Comparison with Previous SR and VSR Approaches
To evaluate the performance of DC-VSR, we compare it with various previous methods across different categories. Specifically, our evaluation includes non-generative image super-resolution (ISR) methods: bicubic interpolation and SwinIR (Liang et al., 2021); two generative ISR methods: Real-ESRGAN (Wang et al., 2021) and StableSR (Wang et al., 2024); two non-generative VSR methods: RealBasicVSR (Chan et al., 2022b) and RealViformer (Zhang and Yao, 2024); and two generative VSR methods: Upscale-A-Video (UAVideo) (Zhou et al., 2024) and MGLD (Yang et al., 2024a).
@@
Table 1 presents a quantitative comparison with previous methods. It is important to note that quantitatively measuring VSR quality is challenging, especially for generative-prior-based approaches, due to the synthesized details that may not align with the ground-truth details. Nevertheless, on both synthetic and real-world datasets, DC-VSR consistently demonstrates superior or comparable results across various metrics. For PSNR and SSIM, non-generative methods generally score higher because they do not synthesize high-frequency details that may not align with ground-truth videos. However, DC-VSR achieves PSNR scores that are competitive and SSIM scores that surpass previous generative VSR methods. When considering DISTS, a perceptual metric, DC-VSR ranks second among all methods. Regarding temporal consistency, measured by tOF and tLP, DC-VSR outperforms all other methods, including non-generative ones, highlighting the effectiveness of video generative prior and our TAP scheme. Finally, evaluating video quality independent of ground-truth data using MUSIQ and DOVER, DC-VSR outperforms other methods in most cases on both synthetic and real-world datasets.
@@
Figure 4. (a)&(b) are VSR results of image diffusion prior-based methods. (c)&(d) show the effects on the proposed TAP.
Fig. 9 presents a qualitative comparison highlighting the restoration quality of a single frame, demonstrating the superiority of DC-VSR over previous methods by a significant margin. For example, DC-VSR excels at restoring building windows, adding plausible details to degraded regions, as seen in VideoLQ 020. Notably, it surpasses other VSR methods in restoring characters, a particularly challenging task for generative models, as seen in VideoLQ 008 and 049. These results indicate that DC-VSR not only ensures spatio-temporal consistency but also achieves outstanding restoration quality.
@@
Figure 5. Visualization of the effects on the proposed SAP.

### v2

@@ Abstract.
Video super-resolution (VSR) aims to reconstruct a high-resolution (HR) video from a low-resolution (LR) counterpart. Achieving successful VSR requires producing realistic HR details and ensuring both spatial and temporal consistency. To restore realistic details, diffusion-based VSR approaches have recently been proposed. However, the inherent randomness of diffusion, combined with their tile-based approach, often leads to spatio-temporal inconsistencies. In this paper, we propose DC-VSR, a novel VSR approach to produce spatially and temporally consistent VSR results with realistic textures. To achieve spatial and temporal consistency, DC-VSR adopts a novel Spatial Attention Propagation (SAP) scheme and a Temporal Attention Propagation (TAP) scheme that propagate information across spatio-temporal tiles based on the self-attention mechanism. To enhance high-frequency details, we also introduce Detail-Suppression Self-Attention Guidance (DSSAG), a novel diffusion guidance scheme. Comprehensive experiments demonstrate that DC-VSR achieves spatially and temporally consistent, high-quality VSR results, outperforming previous approaches. The project page is available at https://daramgc.github.io/docs/Publications/dc-vsr
Video Super-Resolution, Real-World Video, Video Generative Prior
ccs: Computing methodologies Computational photography
Figure 1. Real-world video super-resolution results of our proposed methods and state-of-the-art models: RealViformer (Zhang and Yao, 2024), Upscale-A-Video (Zhou et al., 2024), MGLD (Yang et al., 2024a). DC-VSR shows outstanding super-resolution performance, in respect of not only quality but also spatial and temporal consistency. The input videos are from the VideoLQ dataset (Chan et al., 2022b) (samples 037 and 038).
@@ 1. Introduction
To achieve successful VSR, it is essential to generate realistic HR details while maintaining both spatial and temporal consistency. To this end, over the past few decades, numerous approaches have been proposed, including traditional methods such as interpolation techniques, model-based optimization approaches, and recent neural network-based ones (Zhang and Yao, 2024). Nevertheless, most of them struggle to produce realistic HR details due to a lack of effective priors that model high-frequency details as shown in Fig. 1(b).
Recently, diffusion models that provide powerful generative priors for natural images have been exploited for VSR to achieve high-quality results with realistic textures. Yang et al. (2024a) and Zhou et al. (2024) utilize image diffusion models and successfully restore detail-rich HR videos. However, adopting image diffusion models poses challenges in maintaining spatial and temporal consistency due to the design of image diffusion models, which typically target single images with limited spatial sizes due to significant memory usage. To achieve temporally-consistent VSR using image diffusion models, these approaches adopt additional temporal layers, temporal propagation with motion compensation, and overlapping of temporal windows. Furthermore, to handle large video frames, frames are split into overlapping tiles, processed individually, and then merged. Despite these efforts, resulting videos often suffer from spatial and temporal inconsistencies due to the inherent randomness of image diffusion models and their tile-based approach (Fig. 1(c) and (d)).
In this paper, we propose DC-VSR (Diffusion-based Consistent VSR), a novel VSR approach that, for the first time, leverages a video diffusion prior. Our approach produces spatially and temporally consistent results with realistic textures, given an LR video of arbitrary length and spatial size (Fig. 1(e)). We employ Stable Video Diffusion (SVD) (StabilityAI, 2023) as a video diffusion prior, which provides powerful generative capabilities for restoring high-quality, temporally-consistent details. However, exploiting an existing video diffusion model for consistent VSR over long video clips with large frames is challenging, as existing models are designed to synthesize a limited number of small frames, similar to image diffusion models.
@@
  • We propose SAP that injects subsampled features representing the entire area of a video frame to different tiles, ensuring spatial consistency.
  • We propose TAP for sharing information across temporally distant frames, achieving temporal consistency.
  • We introduce DSSAG to enhance the quality of synthesized video frames without additional computational overhead.
@@ 2. Related Work
VSR methods can broadly be classified into two main categories based on the use of generative prior. VSR models without generative prior (Youk et al., 2024; Chan et al., 2021, 2022a, 2022b; Xie et al., 2023; Zhang and Yao, 2024) are typically trained using reconstruction losses, such as the mean squared error (MSE), augmented with additional techniques designed to enhance temporal consistency. Although they restore HR videos closely matching the input LR videos, they fail to generate complex details, producing blurry results due to the significant ill-posed nature of the problem. To achieve VSR with rich details, generative prior-based VSR methods (Rota et al., 2024; Xu et al., 2024; Chen et al., 2024; Zhou et al., 2024; Yang et al., 2024a) have recently emerged. These methods fine-tune GANs (Goodfellow et al., 2014) or diffusion-based image generation models by modifying them to take LR inputs as conditions and incorporating additional modules to consider temporal consistency. However, maintaining consistency across frames with detailed textures is still challenging (Blattmann et al., 2023a, b), leading to temporal flickering artifacts.
@@
As deep learning techniques have advanced, various image super-resolution (ISR) methods have been proposed. SRCNN (Dong et al., 2015) was the first to introduce a CNN-based SR method. Subsequent studies have applied various network architectures, such as residual networks (Kim et al., 2016a; Ledig et al., 2017), dense networks (Zhang et al., 2018c; Wang et al., 2018b, 2021), Laplacian pyramid networks (Lai et al., 2017), back-projection networks (Haris et al., 2018), recursive structures (Kim et al., 2016b; Tai et al., 2017), and channel attention (Zhang et al., 2018b). While these methods have demonstrated promising SR results, they often struggle to recover high-frequency details. To address this, approaches leveraging the generative prior of pretrained generative models have been introduced exploiting GANs (Chan et al., 2020; Menon et al., 2020) or Diffusion models (Wang et al., 2024; Yu et al., 2024; Lin et al., 2024; Yang et al., 2024b; Wu et al., 2024b; Sun et al., 2024; Wu et al., 2024a). Since the application of these methods to videos leads to severe flickering artifacts due to the absence of temporal consistency, VSR research has been active in addressing this issue.
@@
Diffusion models face two main challenges in generating higher spatial or temporal resolutions: quality degradation from discrepancies between training and inference conditions, and substantial computational costs. To address these challenges, tile-based generation approaches have been proposed (Bar-Tal et al., 2023; Lee et al., 2023; Du et al., 2024; Zhou et al., 2024; Yang et al., 2024a). These methods divide the input into overlapping tiles, perform diffusion sampling on each tile individually, and subsequently merge the tiles. However, their approaches do not consider inter-tile consistency, often generating inconsistent details for the same content across split tiles in an image or distant frames (Zhou et al., 2024; Yang et al., 2024a).
@@ 3. DC-VSR
At each diffusion sampling step $t$, $x_{t}$ and $l$ are first concatenated in an interleaved manner, i.e., $[x_{t,1},l_{1},\cdots,x_{t,N},l_{N}]$, where $x_{t,i}$ is the noisy latent of the $i$-th HR video frame. The concatenated latents are then split into spatio-temporal tiles. We refer to each tile as $x^{\prime}_{t,m,n}$ where $m$ and $n$ are spatial and temporal indices. Each tile is processed through a denoising U-Net, and the processed tiles are merged to obtain the latent representation $x_{t-1}$ at the next sampling step $t-1$. We utilize spatio-temporal tiles of size $64\times 64\times 14$ in the latent space, corresponding to $512\times 512\times 14$ in the image space with a scaling factor of 8. In line with previous approaches (Zhou et al., 2024; Yang et al., 2024a), spatially and temporally neighboring tiles overlap by 50%. Overlapped tiles are blended in the tile merging step in our pipeline using alpha blending with Gaussian weights. To achieve spatial and temporal consistency, DC-VSR alternatingly applies either SAP or TAP at each sampling step until $t$ reaches 00. Finally, $x_{0}$ is fed to a decoder, producing the HR video $I^{\textrm{HR}}$.
DC-VSR employs a tile-based approach to handle lengthy videos with large frames with a video diffusion prior. However, naïvely splitting a video into tiles may introduce spatial and temporal inconsistencies. In image diffusion models like latent diffusion model (LDM) (Rombach et al., 2022), self-attention layers of the denoising U-Nets ensure spatial consistency within an image. Likewise, video diffusion models such as SVD (StabilityAI, 2023) leverage self-attention to achieve spatially and temporally coherent results. Specifically, the self-attention operation is defined as:
@@ 3.3. Detail-Suppression Self-Attention Guidance
SAG and PAG noticeably improve image synthesis quality, especially when combined with CFG. However, integrating them with CFG incurs substantial computational costs. While using $\epsilon_{\theta}(x_{t},c)$ instead of $\epsilon_{\theta}(x_{t})$ in Eqs. 5 and 6 allows this combination, it requires running the denoising U-Net three times for SAG, as it needs additional running of the denoising U-Net to detect high-frequency regions for the blurring operation. For PAG, the fixed level of perturbation complicates balancing the effects of CFG and PAG when combined. Consequently, PAG and CFG are typically applied separately, which also necessitates three runs of the denoising U-Net.
Table 1. Quantitative comparisons of different VSR (x4) models on synthetic datasets (REDS4 and UDM10) and real datasets (VideoLQ). ISR, VSR and GP are image super-resolution tasks, video super-resolution tasks and generative prior. The best and second-best scores are marked in bold and underline. $\text{DOVER}^{*}$ and $\text{tLP}^{*}$ denote DOVER ($\times$100) and tLP ($\times$100) respectively.
  Datasets Metrics ISR w/o GP ISR w/ GP VSR w/o GP VSR w/ GP
@@
Figure 3. Denoised results of unconditional term at the intermediate timestep (t=16 out of 25) (a) without and (b) with DSSAG. The input video is from the VideoLQ dataset (Chan et al., 2022b) (sample 013).
@@ 4. Experiments
To build DC-VSR, we fine-tune Image-to-Video Stable Video Diffusion (I2V-SVD) (StabilityAI, 2023), which adopts the LDM framework (Rombach et al., 2022) with the EDM (Karras et al., 2022) diffusion mechanism. We use the REDS dataset (Nah et al., 2019) to train our model. Following previous work (Chan et al., 2022b; Yang et al., 2024a), we merge 240 training videos and 30 test videos, reorganizing them into 266 training videos and 4 test videos, and refer to the latter as REDS4. We refer the reader to the supplementary material for more implementaion details.
We use the REDS4 (Chan et al., 2022b) and UDM10 (Yi et al., 2019) datasets for VSR evaluation. We construct HR-LR video pairs from the datasets using the real-world degradation pipeline of Chan et al. (2022b), which applies random blur, resizing, noise, JPEG compression, and video compression. Additionally, we use the VideoLQ dataset (Chan et al., 2022b) as a real-world LR dataset.
@@ 4.1. Comparison with Previous SR and VSR Approaches
To evaluate the performance of DC-VSR, we compare it with various previous methods across different categories. Specifically, our evaluation includes non-generative image super-resolution (ISR) methods: bicubic interpolation and SwinIR (Liang et al., 2021); two generative ISR methods: Real-ESRGAN (Wang et al., 2021) and StableSR (Wang et al., 2024); two non-generative VSR methods: RealBasicVSR (Chan et al., 2022b) and RealViformer (Zhang and Yao, 2024); and two generative VSR methods: Upscale-A-Video (UAVideo) (Zhou et al., 2024) and MGLD (Yang et al., 2024a). We follow the default settings from the official repositories of prior works in our experiments.
@@
Figure 4. (a)&(b) are VSR results of image diffusion prior-based methods. (c)&(d) show the effects on the proposed TAP. The input video is from the VideoLQ dataset (Chan et al., 2022b) (sample 008).
Table 1 presents a quantitative comparison with previous methods. It is important to note that quantitatively measuring VSR quality is challenging, especially for generative-prior-based approaches, due to the synthesized details that may not align with the ground-truth details. Nevertheless, on both synthetic and real-world datasets, DC-VSR consistently demonstrates superior or comparable results across various metrics. For PSNR and SSIM, non-generative methods generally achieve higher scores as they do not synthesize high-frequency details that may not align with ground-truth videos. However, this often results in blurry outputs as will be shown in Figs. 7, 8 and 9. Compared to previous generative VSR methods, DC-VSR achieves competitive PSNR scores and higher SSIM scores. When considering DISTS, a perceptual metric, DC-VSR ranks second among all methods. Regarding temporal consistency, measured by tOF and tLP, DC-VSR outperforms all other methods, including non-generative ones, highlighting the effectiveness of video generative prior and our TAP scheme. Finally, evaluating video quality independent of ground-truth data using MUSIQ and DOVER, DC-VSR outperforms other methods in most cases on both synthetic and real-world datasets.
@@
Fig. 9 presents a qualitative comparison highlighting the restoration quality of a single frame, demonstrating the superiority of DC-VSR over previous methods. For example, DC-VSR excels at restoring building windows, adding plausible details to degraded regions, as seen in VideoLQ 020. Notably, it surpasses other VSR methods in restoring characters, a particularly challenging task for generative models, as seen in VideoLQ 049. These results indicate that DC-VSR not only ensures spatio-temporal consistency but also achieves outstanding restoration quality.
@@
Figure 5. Visualization of the effects on the proposed SAP. The input video is from Pexels (©Tima Miroshnichenko).